OpenStudy (anonymous):

the total resistance is 18 kΩ, and the battery's emf is 26.0 V. If the time constant is measured to be 45 µs, calculate each of the following. (a) the total capacitance of the circuit (b) the time it takes for the voltage across the resistor to reach 12.0 V after the switch is closed

OpenStudy (noelgreco):

Well, what do you know about calculating the time constant?

OpenStudy (anonymous):

i already calculated part A which is 2.5 but I'm unsure what to do about part B

OpenStudy (anonymous):

1/RC is the time constant, so C = 1/(1.8E4*4.5E-5) seems like a lot of capacitance, but w/e

OpenStudy (anonymous):

i got .25 is that right

OpenStudy (anonymous):

second part is just solving: \[\huge 12 =26e ^{t*4.5*10^{-5}}\] for t

OpenStudy (anonymous):

how are you getting that? using different number than the ones you posted?

OpenStudy (anonymous):

show me your calculation...

OpenStudy (anonymous):

can you show me step by step im so confused

OpenStudy (anonymous):

C = 1/(1.8E4*4.5E-5)

OpenStudy (anonymous):

there's only one step

OpenStudy (anonymous):

no for part B

OpenStudy (noelgreco):

Marian - .25 what? You can't go with that as a raw number in your calculations. The .25 implies .25 farads - something you're unlikely to see.

OpenStudy (noelgreco):

2.5. sorry. Still you need units, and if you're using micro- or pico- you must say so.

OpenStudy (anonymous):

nvm im still confused

OpenStudy (noelgreco):

I dislike suggesting you're working beyond your coverage, but you may need to go back to the drawing board on this one. Google has many options. Good luck.